OpenStudy (anonymous):

a box measures 4 inches by 6 inches .Find the volume of the box in cubic feet .enter your answer as a fraction .

jhonyy9 (jhonyy9):

you need again one height measure

OpenStudy (anonymous):

would i multiply to find the answer ?

jhonyy9 (jhonyy9):

so if this boxe have a base square so than the volume will be 4*4*6=16*6=96 inches squared

jhonyy9 (jhonyy9):

ok ?

OpenStudy (anonymous):

got it

jhonyy9 (jhonyy9):

good luck bye